# Enhancing LLM Reliability via Explicit Knowledge Boundary Modeling

### v3

@@ 1 Introduction
To address these limitations, we propose a novel alignment objective that enhances model awareness of its knowledge boundaries, enabling models to respond maximally while explicitly distinguishing high/low confidence outputs. We introduce the concepts of ”sure” and ”unsure” to clarify this distinction. A reliable system should deliver near-perfect accuracy for ”sure” predictions and provide helpful information in the “unsure” category, serving as a form of soft rejection, appropriately managing user expectations regarding accuracy.
Building on this concept, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework. EKBM integrates both Fast and Slow systems: a fast-thinking model generates responses annotated with confidence labels, allowing for immediate use of high-confidence outputs, while a slow-thinking refinement model engages in deliberate reasoning to improve low-confidence predictions.
Crucially, EKBM does not merely filter uncertain outputs but leverages them as opportunities for improvement, ensuring both precision and coverage. The collaboration between these two systems strikes a balance between reliability and efficiency, with the self-awareness capability of the fast-thinking model being essential for the system’s effective functioning.
To operationalize this paradigm, we design a training pipeline that aligns the model’s capability boundaries and improves self-awareness. Evaluations on dialogue state tracking (DST) tasks demonstrate that our method effectively enables the model to reliably and accurately distinguish the confidence levels of its output compared to baseline methods. Additionally, we develop an automatic refinement model for ”unsure” outputs. Experiments show that the EKBM system achieves superior overall performance compared to traditional approaches that directly tune models for the task.
@@
We present the EKBM framework, integrating fast-thinking confidence labeling with slow-thinking refinement, along with metrics for evaluating LLM self-awareness and reliability.
@@ 2.1 Alignment for Reliability
where denotes a scoring function, is the input prompt, and and denote correct and incorrect responses, respectively, ensuring the score of correct responses exceeds that of incorrect ones.
@@
While this rejection strategy aligns with reliability goals, it may unduly limit the model’s helpfulness in practical scenarios. Therefore, we propose a more nuanced objective: the model should aim to provide answers whenever feasible, categorizing outputs into high-confidence (”sure”) and low-confidence (”unsure”) responses. The model must ensure high accuracy in the ”sure” category, while tolerating some errors in the ”unsure” category, which can serve as references for users or inputs for further refinement. Our proposed alignment objective is formalized as:
Here, and represent correct and incorrect responses, respectively, while and represent ”sure” and ”unsure” as levels of confidence. This objective guarantees that correct ”sure” predictions are prioritized, followed by correct ”unsure” predictions, which, while less confident, still provide utility. Incorrect predictions in the ”unsure” category are discouraged but deemed more acceptable than those in the ”sure” category, as the latter significantly undermine overall reliability.
@@ 2.2.1 Weighted-F1
F1-score is usually taken as the evaluation metric for multi-slot problems:
where TP, FP and FN represent True Positives, False Positives and False Negatives respectively.
We modify the F1 score to align with the objectives of our EKBM framework. Since ”unsure” predictions represent an state of uncertainty, we assign a weighted contribution (denoted by and , both ranging from 0 to 1) to reflect their partial utility as candidates for refinement. The modified precision and recall are defined as:
where STP and SFP denote Sure True Positives and Sure False Positives, while UTP and UFP represent the corresponding cases in the Unsure category. For UTP in the recall calculation, we only require that a prediction be partially correct (i.e., refinable). Unlike precision, we do not introduce the weighting parameter in the denominator of recall. This is because recall measures the proportion of correct predictions relative to the total number of gold labels, which remains unchanged regardless of the introduction of . Notably, incorporating does not affect the range of precision and recall, both of which remain within [0,1]. The modified Weighted-F1 score is defined as:
@@ 2.2.2 Metric-Objective Consistency: Weighted-F1 as a Parametric Alignment Measure
The Weighted-F1 metric structurally aligns with our four-tiered alignment objective (Equation (3)) through a parametric formulation that encodes hierarchical priorities. Its design ensures: (1) Strict enforcement for sure predictions: Unweighted sure terms prioritize high-confidence correctness (STP) while imposing maximal penalties for sure errors (SFP); (2) Controlled utility for unsure predictions: -scaled unsure terms modulate low-confidence outputs—attenuated UFP penalties tolerate unsure errors, whereas partial UTP rewards incentivize refinable correct predictions.
@@
We discuss a few corner cases: (1) = 0, = 0: Weighted-F1 disregards all ”unsure” predictions. (2) = 0, = 1: Errors in the ”unsure” category are ignored, rewarding effective recall, representing the theoretical upper limit after perfect refinement. (3) = 1, = 0: Lacks significant interpretation. (4) = 1, = 1: Eliminates confidence labeling, reverting to traditional methods.
@@ 2.2.3 Reliability Metrics
We employ two specific Weighted-F1 scores as performance metrics. Among these, Quality-F1 is the core metric for assessing model reliability, while Optimal-F1 serving as a supplementary reference.
Quality-F1 Defined as , it assigns ”unsure” predictions half the weight of ”sure” ones, encouraging the model to select ”sure” when confident while minimizing penalties for ”unsure” outputs. Quality-F1 measures the model’s self-awareness ability and reliability.
Optimal-F1 Defined as , it represents the theoretical upper limit achievable after perfect refinement of ”unsure” predictions, reflecting the overall potential of the two-stage system.
@@ 3.1 Explicit Knowledge Boundary Modeling
We formalize our proposed framework, Explicit Knowledge Boundary Modeling (EKBM), which aims to improve the reliability and self-awareness of the model by explicitly modeling its knowledge boundaries and improve overall performance by incorporating a refinement model to improve the ”unsure” predictions. As shown in Figure 2 (1), the framework divides the model’s decision-making process into two distinct stages, balancing immediate usability with comprehensive coverage, making the system both reliable and helpful.
@@
In the first stage, the model performs the primary task while simultaneously assigning a confidence label (”sure” or ”unsure”) to each prediction. Predictions labeled as ”sure” are directly accepted as final outputs, ensuring high precision.
@@
In the second stage, ”unsure” predictions undergo further refinement through strategies like user confirmation, automated multi-step reasoning, or other post-processing techniques, incuring extra cost but increases overall accuracy and reliability of the system.
@@ 3.2 Reliability Alignment
To align the model’s behavior with our proposed alignment objective, we explore various training strategies to enhance its intrinsic self-awareness (Xu et al., 2025). We apply Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) for alignment, with the data construction process illustrated in Figure 2(2) and two detailed examples provided in Appendix §B.
@@
We assume that the model’s task-specific knowledge boundary can be approximated by repeatedly sampling on the training data. Accordingly, our SFT data construction approach is straightforward: we conduct multiple sampling for each training sample. Predictions consistently deemed accurate are labeled ”sure,” while those with errors or omissions are marked ”unsure.” Crucially, we intentionally refrain from correcting erroneous outputs to avoid introducing supervisory signals that might inadvertently improve the model’s task performance. That is, when the model produces an incorrect prediction, we simply assign a confidence label of ”unsure” without rectifying the error. The erroneous output remains in the final training data, reinforcing the model’s ability to recognize and categorize uncertain information rather than improving its accuracy directly.
The number of sampling rounds, denoted as i, is a hyperparameter. A larger i leads to a more conservative model, as the proportion of ”unsure” labels increases with greater i in the training corpus. In our study, we set i = 1 by default to accurately reflect the model’s realistic capability boundary.
@@ 3.3 Refinement Model
We train automated refinement models for each dataset to further optimize unsure predictions by multi-step reasoning, leveraging supervised fine-tuning (SFT) based on the LLAMA3 8B model. Our refinement operates at a fine-grained level, addressing each unsure prediction individually. We employ a reasoning paradigm similar to the Chain of Thought (Wei et al., 2022) methodology and use GPT-4o (Hurst et al., 2024) to generate detailed reasoning process for each training prediction. This approach is illustrated in Figure 2 (3), and detailed information can be found in Appendix §C.
@@ 4 Experiments
RQ1 Does our reliability training pipeline outperform current methods in enhancing self-awareness?
RQ2 Does our EKBM framework improve the overall performance of LLMs on complex tasks?
@@ 4.1.1 Model and Baselines
We utilize the LLaMA3 8B model and the Qwen2.5 7B model as the backbone for our experiments.
For reliability alignment, we conducted Reliability-SFT and integrated DPO to create two additional variants: Reliability-SFT+DPO Joint Training, which undergoes joint training with SFT and DPO, and Reliability-SFT+DPO Post Training, where DPO training follows initial SFT training, optimizing the model beyond the Reliability-SFT foundation. The key distinction lies in the different source distributions of preference data sampled for DPO. Training details could be found in Appendix §A.
To evaluate the reliability of our proposed models, we compare them against several baselines, including three prompt-based and three uncertainty-based approaches (see Appendix §D). Notably, our focus is on the model’s self-awareness rather than task performance, so we do not include SOTA DST models for comparison. By default, we utilize only 1,000 samples for Reliability-SFT training and 2,000 for DPO training—significantly less than traditional models.
@@
methods encompass three approaches: Direct, Verbose, and Self-Reflection (SR). The Direct approach generates predictions without confidence estimation, whereas Verbose simultaneously produces confidence labels. SR assesses each prediction through self-evaluative reflection.
@@
methods include Token Probability (Prob), Self-Consistency (SC), and P(True) (Kadavath et al., 2022), each employing distinct techniques to classify outputs as ”sure” or ”unsure”. Prob calculates average token probabilities, SC measures the frequency of predictions across multiple samples, and P(True) determines the probability of the ”True” token during model self-evaluation.
Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA3-8B Dubey et al. (2024) Prompt Direct 72.48 77.98 73.38 73.38 73.37 64.83 65.39 65.39 26.46 28.71 25.12 25.12 Verbose 85.95 85.85 80.88 53.34 91.53 76.52 78.09 53.56 69.59 27.93 29.25 14.98 SR 86.43 81.34 81.34 61.43 92.39 76.32 78.84 48.60 65.32 33.61 33.37 21.30 Uncertainty Prob 80.05 80.66 77.74 69.03 83.50 72.44 73.40 60.30 36.93 32.28 27.74 22.86 SC 80.11 89.51 84.15 64.48 83.16 75.48 77.33 62.93 84.66 41.06 42.76 16.55 P(True) 72.82 79.00 73.50 70.81 72.44 66.26 65.97 63.06 24.78 29.93 26.26 26.24 Reliability (Ours) SFT 94.04 91.75 91.65 82.44 95.66 88.25 88.56 83.08 86.74 85.51 82.66 61.64 + DPO Joint 94.52 92.16 92.83 83.50 95.84 88.85 88.89 83.11 89.07 86.31 85.89 61.42 + DPO Post 93.59 93.10 91.75 83.55 95.17 88.00 88.35 85.78 82.26 83.47 79.70 64.14 Qwen2.5-7B Yang et al. (2024) Prompt Direct 69.42 71.37 67.22 67.22 71.19 65.31 65.98 65.98 33.43 36.11 33.40 33.40 Verbose 83.02 71.14 69.05 58.88 78.62 77.22 75.23 64.78 65.78 31.80 32.13 21.74 SR 73.54 72.08 69.69 67.87 73.73 66.48 67.90 66.34 35.87 36.82 35.13 34.05 Uncertainty Prob 77.12 74.47 72.02 65.19 84.57 74.43 75.49 61.11 46.66 38.91 34.83 29.36 SC 83.33 90.20 84.67 57.00 87.65 74.83 77.42 51.43 71.47 61.08 53.38 25.22 P(True) 70.03 71.44 67.49 67.24 72.11 65.66 66.13 65.65 33.60 36.40 33.67 33.56 Reliability (Ours) SFT 94.16 91.93 91.98 82.03 96.08 89.13 89.33 83.01 90.69 83.86 83.09 64.63 + DPO Joint 95.29 92.88 93.03 80.86 95.99 89.86 89.51 82.00 91.23 85.78 83.95 64.48 + DPO Post 94.02 92.23 92.12 82.89 96.15 89.29 89.60 84.12 86.70 84.61 81.00 67.51
@@ 4.1.2 Dataset
We evaluate our methods on the Dialogue State Tracking (DST) task in task-oriented dialogue systems, which involves extracting slot-value pairs from multi-turn dialogues based on a predefined ontology (Xu et al., 2024a). We focus on the intricate multi-slot problem, where the model is likely to exhibit varying confidence levels for different parts of responses to the same query. Consequently, we do not adopt traditional QA or mathematical datasets. To ensure a robust evaluation, we utilize three datasets: MultiWOZ-2.4 (Ye et al., 2021), BiTOD (Lin et al., 2021) and SGD (Rastogi et al., 2020). For BiTOD, we adopt the English version, while for SGD, we randomly select 10,000 instances from the testset to mitigate excessive computational overhead.
Common evaluation metrics for DST include Joint Goal Accuracy (JGA) and Slot-F1. For JGA, a sample is deemed correct only if all predictions are accurate without omissions, while Slot-F1 offers a slot-level assessment.

### v4

@@ 1 Introduction
To address these limitations, we propose a novel alignment objective that enhances model awareness of its knowledge boundaries, enabling models to respond more effectively by explicitly distinguishing between high-confidence (“sure”) and low-confidence (“unsure”) outputs. A reliable system should deliver near-perfect accuracy for “sure” predictions and provide helpful information in the “unsure” category, serving as a form of soft rejection, appropriately managing user expectations regarding accuracy. Building on this concept, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework. EKBM integrates both Fast and Slow systems: a fast-thinking model generates responses annotated with confidence labels, allowing for immediate use of high-confidence outputs, while a slow-thinking refinement model engages in deliberate reasoning to improve low-confidence predictions. Crucially, EKBM does not merely filter uncertain outputs but
leverages them as opportunities for improvement, ensuring both precision and recall. The collaboration between these two systems strikes a balance between reliability and efficiency, with the fast-thinking model’s self-awareness being crucial for the system’s effective functioning.
To operationalize this paradigm, we design a training pipeline that aligns the model’s capability boundaries and improves self-awareness. Evaluations on dialogue state tracking (DST) tasks demonstrate that our method effectively and reliably distinguishes the confidence levels of model output compared to baseline methods. Additionally, we develop an automatic refinement model for “unsure” outputs. Experiments show that the EKBM system achieves superior overall performance compared to traditional approaches that directly tune models for the task.
@@
We present the EKBM framework, integrating fast-thinking assessment with slow-thinking refinement, along with metrics for evaluating self-awareness and reliability.
@@ 2.1 Alignment for Reliability
where a scoring function ensures that the score of a correct response to an input prompt is higher than that of an incorrect response .
@@
While this rejection strategy ensures reliability, it may unduly limit the model’s helpfulness. Therefore, we propose a more nuanced objective: the model should provide an answer whenever possible, categorizing its output as either high-confidence (“sure”) or low-confidence (“unsure”). The model must ensure high accuracy for “sure” responses while tolerating some errors in the “unsure” category, which can serve as references for users or inputs for further refinement. Our proposed alignment objective is formalized as:
Here, and denote correct and incorrect responses, while and represent “sure” and “unsure” confidence levels. This objective guarantees that correct “sure” predictions are prioritized over correct “unsure” predictions, which, while less confident, still provide utility. Incorrect “unsure” predictions are discouraged but considered more acceptable than incorrect “sure” predictions, as the latter significantly undermine overall reliability.
@@ 2.2.1 Weighted-F1
The F1-score, calculated using True Positives (TP), False Positives (FP), and False Negatives (FN), is typically the evaluation metric for multi-slot problems:
We modify the F1 score to align with the objectives of our EKBM framework. Since “unsure” predictions represent an state of uncertainty, we assign a weighted contribution (denoted by and , both ranging from 0 to 1) to reflect their partial utility as candidates for refinement. The modified precision and recall are defined as:
Here, STP and SFP denote True Positives and False Positives for “sure” predictions, while UTP and UFP are the corresponding values for the “unsure” category. For recall, we consider a prediction to be a UTP if it is partially correct (i.e., refinable). Unlike precision, we do not apply the weighting parameter to the denominator of the recall calculation, as recall measures the proportion of gold labels retrieved, which is independent of the introduction of . Notably, incorporating does not affect the range of precision and recall, both of which remain within [0,1]. The modified Weighted-F1 score is defined as:
@@ 2.2.2 Metric-Objective Consistency: Weighted-F1 as a Parametric Alignment Measure
The Weighted-F1 metric aligns with our four-tiered objective (Equation (3)) by using a parametric formulation that encodes hierarchical priorities, which ensures: (1) Strict enforcement for sure predictions: Unweighted terms prioritize high-confidence correctness (STP) and apply maximal penalties for errors (SFP). (2) Controlled utility for unsure predictions: The -scaled terms modulate low-confidence outputs, with attenuated UFP penalties tolerating unsure errors, and partial UTP rewards incentivizing refinable correct predictions.
@@
We discuss a few corner cases: (1) = 0, = 0: Weighted-F1 disregards all “unsure” predictions. (2) = 0, = 1: Errors in the “unsure” category are ignored, rewarding effective recall, representing the theoretical upper limit after perfect refinement. (3) = 1, = 0: Lacks significant interpretation. (4) = 1, = 1: Eliminates confidence labeling, reverting to traditional methods.
@@ 2.2.3 Reliability Metrics
We use two specific Weighted-F1 scores: Quality-F1 as the primary metric for assessing model reliability, and Optimal-F1 as a supplementary reference.
Quality-F1: Defined as , this metric assigns half the weight to “unsure” predictions compared to “sure” ones. This encourages the model to provide high-confidence “sure” outputs while mitigating penalties for “unsure” ones. Quality-F1 directly measures the model’s self-awareness and reliability.
Optimal-F1 Defined as , it represents the theoretical upper limit achievable after perfect refinement of “unsure” predictions, reflecting the overall potential of the system.
@@ 3.1 Explicit Knowledge Boundary Modeling
We formalize our proposed framework, Explicit Knowledge Boundary Modeling (EKBM), which aims to improve model reliability and self-awareness by explicitly modeling its knowledge boundaries and improve overall performance by incorporating a refinement model to improve the “unsure” predictions. As shown in Figure 2 (1), the framework divides the model’s decision-making process into two distinct stages, balancing immediate usability with comprehensive coverage, making the system both reliable and helpful. We provide a complete example of the two-stage execution flow and model outputs in Appendix B.
@@
In the first stage, the model performs the primary task and simultaneously assigns a confidence label (“sure” or “unsure”) to each prediction. “Sure” predictions are directly accepted as final outputs, ensuring high precision.
@@
In the second stage, “unsure” predictions undergo further refinement through strategies like user confirmation, automated multi-step reasoning, or other post-processing techniques, incuring extra cost but increases overall accuracy and reliability of the system.
@@ 3.2 Reliability Alignment
To align the model’s behavior with our alignment objective, we explore various training strategies to enhance its intrinsic self-awareness (Xu et al., 2025), including Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO). The data construction process is shown in Figure 2(2), with two detailed examples provided in Appendix §C.
@@
We assume that a model’s task-specific knowledge boundary can be approximated by repeated sampling on training data. Accordingly, our SFT data construction is straightforward: for each training sample, we perform multiple samplings. Predictions consistently deemed accurate are labeled “sure,” while those with errors or omissions are marked “unsure.” Crucially, we intentionally refrain from correcting erroneous outputs to avoid introducing supervisory signals that might inadvertently improve the model’s task performance. That is, when the model produces an incorrect prediction, we simply assign a confidence label of “unsure” without rectifying the error. The erroneous output remains in the final training data, reinforcing the model’s ability to recognize and categorize uncertain information rather than improving its accuracy directly.
The number of sampling rounds i is a hyperparameter. A larger i results in a more conservative model, as the proportion of “unsure” labels in the training corpus increases. In our study, we set i = 1 by default to accurately reflect the model’s realistic capability boundary.
@@ 3.3 Refinement Model
We train automated refinement models for each dataset to further optimize unsure predictions by multi-step reasoning, leveraging supervised fine-tuning (SFT) based on the LLAMA3 8B model. Our refinement operates at a fine-grained level, addressing each unsure prediction individually. We employ a reasoning paradigm similar to the Chain of Thought (Wei et al., 2022) methodology and use GPT-4o (Hurst et al., 2024) to generate detailed reasoning process for each training prediction. This approach is illustrated in Figure 2 (3), and detailed information can be found in Appendix §D.
@@ 4 Experiments
RQ1 Does our training pipeline outperform current methods in enhancing self-awareness?
RQ2 Does EKBM framework improve the overall performance of LLMs on complex tasks?
@@ 4.1.1 Model and Baselines
We utilize the LLaMA3 8B model and the Qwen2.5 7B model as the backbone for our experiments. For reliability alignment, we conducted Reliability-SFT and integrated DPO to create two additional variants: Reliability-SFT+DPO Joint Training, which undergoes joint training with SFT and DPO, and Reliability-SFT+DPO Post Training, where DPO training follows initial SFT training, optimizing the model beyond the Reliability-SFT foundation. The key distinction lies in the different source distributions of preference data sampled for DPO. Training details could be found in Appendix §A.
To evaluate the reliability of our proposed models, we compare them against six baselines: three prompt-based and three uncertainty-based approaches (see Appendix §E). Notably, our focus is on the model’s self-awareness rather than task performance, so we do not include SOTA DST models. By default, we use only 1,000 samples for Reliability-SFT training and 2,000 for DPO, which is significantly less than traditional models.
@@
methods encompass three approaches: Direct, Verbose, and Self-Reflection (denoted as SR, Ji et al. (2023a)). The Direct approach generates predictions without confidence estimation, while Verbose produces both predictions and confidence labels simultaneously. SR assesses each prediction through self-evaluative reflection.
@@
methods include Token Probability (Prob, Manakul et al. (2023)), Self-Consistency (SC, Chen & Mueller (2024)), and P(True) (Kadavath et al., 2022). These methods use different techniques to classify outputs as “sure” or “unsure”. Prob calculates average token probabilities; SC measures the frequency of predictions across multiple samples; and P(True) determines the probability of the “True” token during model self-evaluation.
Method Type Method MultiWOZ-2.4 BiTOD SGD Opti. F1 Qual. F1 Opti. F1 Qual. F1 Opti. F1 Qual. F1 LLAMA3-8B (Dubey et al., 2024) Prompt Direct 72.48 77.98 73.38 73.38 73.37 64.83 65.39 65.39 26.46 28.71 25.12 25.12 Verbose 85.95 85.85 80.88 53.34 91.53 76.52 78.09 53.56 69.59 27.93 29.25 14.98 SR 86.43 81.34 81.34 61.43 92.39 76.32 78.84 48.60 65.32 33.61 33.37 21.30 Uncertainty Prob 80.05 80.66 77.74 69.03 83.50 72.44 73.40 60.30 36.93 32.28 27.74 22.86 SC 80.11 89.51 84.15 64.48 83.16 75.48 77.33 62.93 84.66 41.06 42.76 16.55 P(True) 72.82 79.00 73.50 70.81 72.44 66.26 65.97 63.06 24.78 29.93 26.26 26.24 Reliability (Ours) SFT 94.04 91.75 91.65 82.44 95.66 88.25 88.56 83.08 86.74 85.51 82.66 61.64 + DPO Joint 94.52 92.16 92.83 83.50 95.84 88.85 88.89 83.11 89.07 86.31 85.89 61.42 + DPO Post 93.59 93.10 91.75 83.55 95.17 88.00 88.35 85.78 82.26 83.47 79.70 64.14 Qwen2.5-7B (Yang et al., 2024) Prompt Direct 69.42 71.37 67.22 67.22 71.19 65.31 65.98 65.98 33.43 36.11 33.40 33.40 Verbose 83.02 71.14 69.05 58.88 78.62 77.22 75.23 64.78 65.78 31.80 32.13 21.74 SR 73.54 72.08 69.69 67.87 73.73 66.48 67.90 66.34 35.87 36.82 35.13 34.05 Uncertainty Prob 77.12 74.47 72.02 65.19 84.57 74.43 75.49 61.11 46.66 38.91 34.83 29.36 SC 83.33 90.20 84.67 57.00 87.65 74.83 77.42 51.43 71.47 61.08 53.38 25.22 P(True) 70.03 71.44 67.49 67.24 72.11 65.66 66.13 65.65 33.60 36.40 33.67 33.56 Reliability (Ours) SFT 94.16 91.93 91.98 82.03 96.08 89.13 89.33 83.01 90.69 83.86 83.09 64.63 + DPO Joint 95.29 92.88 93.03 80.86 95.99 89.86 89.51 82.00 91.23 85.78 83.95 64.48 + DPO Post 94.02 92.23 92.12 82.89 96.15 89.29 89.60 84.12 86.70 84.61 81.00 67.51
@@ 4.1.2 Dataset
We evaluate our methods on the Dialogue State Tracking (DST) task in task-oriented dialogue systems, which involves extracting slot-value pairs from multi-turn dialogues based on a predefined ontology (Xu et al., 2024a). We focus on the intricate multi-slot problem, where the model is likely to exhibit varying confidence levels for different parts of responses to the same query. Consequently, we do not adopt traditional QA or mathematical datasets. To ensure a robust evaluation, we utilize three datasets: MultiWOZ-2.4 (Ye et al., 2021), BiTOD (Lin et al., 2021) and SGD (Rastogi et al., 2020). For BiTOD, we adopt the English version, while for SGD, we randomly select 10,000 instances from the testset to mitigate excessive computational overhead. Common evaluation metrics for DST include Joint Goal Accuracy (JGA) and Slot-F1. For JGA, a sample is deemed correct only if all predictions are accurate without omissions, while Slot-F1 offers a slot-level assessment.